# Improved Contact Graph Routing in Delay Tolerant Networks with Capacity and Buffer Constraints

### v1

@@ II-B4 Additional comments on CGR
Although widely used, the CGR is computationally complex. Each node has to perform many checks for each bundle to only determine the next contact to use. Moreover, since these checks are performed after the route search, valid solutions may not be found among the candidate routes. This may trigger a time-consuming iterative search process between the Dijkstra algorithm and the checks.
II-C Source-routing CGR
@@
One solution to reduce CGR’s complexity is Source Routing (SR) [15]. Unlike per hop routing, the idea of SR is to compute the route at the source node and encode it into the bundle in an extension block [16].
@@
When the bundle, whose route has already been calculated, arrives at an intermediate node, this node only checks the validity of the encoded route. In case of invalidity, traditional CGR is performed from this intermediate node. The following two types of SR can be implemented.
II-C1 Bundle-independent SR
@@
With bundle-independent SR, a source performs the route search without bundle properties and queuing information. During the forwarding time, the checks of CGR such as ETO and volume availability are performed. The validity of the encapsulated route is checked using available information on bundles, like its size, and local queue states [17].
II-C2 Bundle-dependent SR
@@
When available at the initial route search time, the bundle properties, e.g., size and priority, can be utilized [5], as is the case with bundle-dependent SR. Then, during the forwarding phase, information on local queues is taken into account and the viability of the route is checked.
@@
Our solution goes further in this paradigm by incorporating network information alongside bundle properties during the initial route search.
@@ III-A Contact-capacity management
As mentioned previously, in CGR capacity consumption and availability are managed during the forwarding phase. The common approach to take contact capacity into account in the route search is to check that the available volume of contacts is greater than the size of the bundle. This is the so called linear approach [5]. However, this does not prevent exceeding capacity limits: It does not track the time at which the bundle uses the contact. The linear volume model is represented on the right side of Figure 3, see also [5].
We detail this issue with the example of Figure 3. In this example, when using the linear model, one may think that bundle $i$ (hatched rectangle) reaching node C at $t_{i}$ can use contact $C_{3}$ immediately. Nevertheless, in reality the contact volume is consumed as represented on the left side of Figure 3. In this case, a bundle reaching node C at $t_{i}$ cannot be transmitted using contact $\text{C}_{3}$ immediately because bundle $a$ (striped rectangle) is being transmitted and uses all the capacity at this specific time. In traditional CGR, this problem is detected during the forwarding step, after the route search, using the ETO for example. The transmission of the bundle should thus be delayed accordingly.
@@
Figure 2: Initial CP.
@@ III-B Buffer management
In DTNs, node buffer management is an important task, as bundles can wait a long time in node buffers before being transmitted. In addition, available buffer capacities may be limited.
Buffer management strategies usually consist of addressing a problem once it has occurred. For example when a buffer is full, the decision to reroute or drop bundles takes place on the spot, like in reactive custody transfer [2]. Authors in [18] suggested a strategy to refuse custody in order to preserve buffer space at nodes. However, these strategies rely only on information coming from local neighbor nodes.
An alternative approach would be to have a global vision on buffer occupancy and to handle the buffer space issues during the initial route search. Note however that, unlike the volume and capacity that are managed at the contact level, the buffer occupation is managed at the node level. For instance, multiple bundles using distinct contacts can affect the buffer of the same node. Therefore, one should be careful when addressing this challenge in a contact-graph representation of the network.
@@ IV Proposed global capacity and buffer management solution
The framework considered is the bundle-dependent SR-CGR [15], see Section II-C2. We propose to address the buffer and capacity constraints before or during the route search. As a reminder, this is to be opposed with approaches reported in the state-of-the-art where actions are taken once an issue has occurred.
The solution ensures that any route found respects the buffer and capacity limits. Moreover, the proposed modifications do not remove any viable solution, which enables the route search to find the optimal solution given the constraints.
This is achieved by modifying the CP before starting the route search, both for the capacity and buffer constraints. Moreover, for the buffer constraints, the allowed edges between contacts are also modified during a route search. In other words, we modify the parameters such that CGR performs the search only among the valid solutions.
The optimal route that respects buffer and capacity limits is then encoded in the bundle, which is forwarded through the network. This manner of sharing the bundle transmission path is the same as in source-routing CGR.
@@ IV-A Model-parameter modifications
  • First, we propose to modify the CP to be used by a route-search algorithm (e.g., Dijkstra search) via contact splitting. We differentiate between two kinds of CP modifications: The definitive ones, which are to be shared in the network, and the temporary bundle-specific modifications, which are performed locally by one node within its route-search process for a specific bundle.
  • Second, we propose imposing additional restrictions on allowed edges between contacts. These rules are then to be respected inside the route-search algorithm. These restrictions are to be implemented only locally by a node, during its route search for a specific bundle.
@@ IV-A1 Contact splitting to modify the CP
We introduce the contact splitting operation. Let us consider an original arbitrary contact $\text{C}_{j}$, having a data rate $R_{C_{j}}$, between a sender node $M$ and a receiver node $N$ from $t_{1,j}$ to $t_{2,j}$. Contact splitting consists in replacing the original contact in the CP, by two split contacts $\text{C}_{j,1}$, $\text{C}_{j,2}$ between nodes $M$ and $N$: one from $t_{1,j}$ to $t_{e_{1}}$ and the other one from $t_{e_{2}}$ to $t_{2,j}$, such that $t_{1,j}\leq t_{e_{1}}\leq t_{e_{2}}\leq t_{2,j}$, see Figure 4. Split contacts have the same data rate $R_{C_{j}}$ as the original contact $C_{j}$.
Note that if $t_{e_{2}}-t_{e_{1}}>0$, part of the original contact is erased. Moreover, one of the split contacts might last for no time, meaning that contact splitting results in contact shortening.
@@ IV-A2 Edge pruning
We define edge pruning as the act of forbidding edges between contacts. This action can be taken between any pair of contacts (original or split). For example, in Figure 3, one can choose to forbid the edge between contacts $\text{C}_{1}$ and $\text{C}_{2}$. This means that, during a route search, there is no option to choose a route having contact $\text{C}_{1}$ followed by contact $\text{C}_{2}$.
@@ IV-B1 Capacity management
We assume knowledge of the bundle size. First, the route search for the first transmitted bundle $i$ is performed. A contact is skipped if the bundle size is greater than the contact volume (similarly to the linear-volume approach). Once the shortest path is selected for this bundle, the CP is modified as follows: We remove the part of the contact consumed by the bundle from the CP before performing the route search for subsequent bundles. The length of the contact to be erased is $t_{\text{Tx,}i}^{C_{j}}$, starting from the start time of the contact or the arrival time of the bundle at the contact’s sender node, whichever occurs later. This variable $t_{\text{Tx,}i}^{C_{j}}$ represents the time needed to transmit bundle $i$ over contact $C_{j}$ in the selected path:
$t_{\text{Tx,}i}^{C_{j}}=\frac{S_{i}}{R_{C_{j}}}.$ (1)
This operation is a contact splitting operation as it creates two new contacts. These two new contacts are to be added to the CP and the original contact is to be removed. This operation is performed for all contacts in the selected route. Then, the subsequent route searches use this modified CP.
The CP modification, to be performed after each selection of a route for a bundle, is definitive and has to be shared with all nodes performing route searches.
Example. In the example of Figure 3, after route search and route selection for the initial bundle $a$ (striped rectangle), the CP becomes as represented in Figure 5. The new CP is then used for route search of subsequent bundles. In this case, the part of contact $\text{C}_{3}$ between $t_{i}$ and $t_{i}+t_{\text{Tx,a}}^{C_{3}}$ is not an option for future bundle and no collision will therefore occur.
Figure 5: Modified CP after allocating route for bundle $a$.
@@ IV-B2 Buffer management
We first introduce the notion of forecast buffer tables. The goal of these tables is to store the buffer occupancy status of each node as a function of time, including future time. Similarly to the CP, we assume that these tables are shared by all entities performing route searches. We then explain how to use these tables to respect the buffer constraints.
Forecast buffer tables. We assume that each node has a maximum buffer capacity limit $B_{\text{max}}$ and that an overflow occurs if this limit is exceeded.
The proposed buffer management solution uses temporary bundle-specific CP modifications, not shared in the network (see the following subsection). However, once a route is selected, the forecast buffer tables are updated accordingly and shared to all entities performing route search. In other words, once a route is validated, one checks when the bundle will reach a given node and how long it will stay in its buffer according to the validated route.
@@
Algorithm to respect the buffer constrains. The main ideas of the proposed algorithm to enforce buffer-limit constraints are as follows. Since the bundle size and the forecast buffer tables are available at the route-search time, one can check whether the bundle can be in the buffer of any node at any time. If not, the algorithm must forbid routes leading to the storage of the bundle in potentially problematic buffers. Hence, before performing a route search, the CP is temporarily modified using contact splitting. Then, some edges are pruned using edge pruning inside the route-search algorithm. This algorithm can be formulated with the four following main steps.
@@
Example. Some problematic contacts for the potential overflow $o_{1}$ descried in the previous example are shown in Figure 8. These contacts have node $Z_{1}$ as receiver (regardless of their sender node $X$), start before $t_{2,o_{1}}$, and end after $t_{1,o_{1}}$. They are problematic because their duration intersects the period of overflow $o_{1}$. The problematic portions of these contacts are horizontally striped in Figure 8. They may forward bundle b to $Z_{1}$ during the potential overflow $o_{1}$ duration. The parts of these contacts that intersect the overflow duration are to be temporarily removed using contact splitting. The same applies to all potential overflows at all nodes.
@@
Example. After step 2, there is no contact having $Z_{1}$ as receiver that spans the duration of $o_{1}$ in Figure 8. Nevertheless, there are contacts having $Z_{1}$ as receiver and ending before $t_{1,o_{1}}$. Using one of these contacts (e.g., $C_{i}$), a bundle arrives to $Z_{1}$ before the potential overflow $o_{1}$. However, this bundle must leave $Z_{1}$ before the beginning of overflow $o_{1}$ to avoid the potential problem. The bundle leaves node $Z_{1}$ through a contact succeeding the previous one ($C_{i}$) in a route. The succeeding contact (e.g., $C_{j}$) has $Z_{1}$ as sender, regardless of the receiver. As a result, the succeeding contact $C_{j}$ should start before overflow $o_{1}$ to avoid it. Otherwise, the succession of $C_{i}$ and $C_{j}$ in a route causes the overflow and is therefore problematic.
@@
Step 4: Finally, the last step is to forbid edges between contacts yielding problematic contact successions. In other words, the edge with the contacts starting too late are pruned. This edge pruning is performed during the route search.
Example. In the example of Figure 8, during the route search, if a contact $C_{i}$ forwards the bundle to node $Z_{1}$ before $o_{1}$, the edge between $C_{i}$ and a contact $C_{j}$ from $Z_{1}$ to a third node is only accepted if $C_{j}$ starts before $t_{1,o_{1}}$. The same applies to all overflows at all nodes.
@@
As described by the above steps, only invalid routes are removed by the modifications of the proposed contact capacity and buffer management solutions. In other words, the proposed modifications before and during the route search do not omit any feasible solution. Hence, the subsequent use of CGR gives the optimal route in the subset of valid solutions.

### v2

@@ II-C4 Additional comments on CGR
Although widely used, CGR is computationally complex. Each node must perform multiple checks per bundle simply to determine the next contact. Since validation occurs after the route search, the system might find no valid solutions among the candidate routes. Such situations often trigger time-consuming iterations between Dijkstra’s algorithm and validation checks.
II-D Source-routing CGR
@@
Source Routing (SR) [31] offers an effective approach to reduce CGR’s complexity. Unlike per-hop routing, SR computes the complete route at the source node and encodes it into the bundle’s extension block [32].
@@
When a bundle with a pre-calculated route arrives at an intermediate node, a verification of the encoded route’s validity is performed. Upon detecting an invalid route, the intermediate node reverts to traditional CGR to compute a new route to the destination. The following two types of SR can be implemented.
II-D1 Bundle-independent SR
@@
In bundle-independent SR, the source node computes routes without considering bundle properties or queuing information. During forwarding, the system performs standard CGR checks such as ETO and volume availability. The system validates the encapsulated route using bundle properties (e.g., size) and local queue states [33].
II-D2 Bundle-dependent SR
@@
Bundle-dependent SR utilizes available bundle properties (e.g., size and priority) during initial route computation [13]. During forwarding, the system evaluates route viability using local queue information.
@@
Our solution extends this approach by incorporating both network information and bundle properties during initial route computation.
@@ III-A Contact-capacity management
CGR manages capacity consumption and availability during the forwarding phase. The route search implements the so-called linear approach that verifies whether available contact volume exceeds the bundle size [13]. This approach fails to prevent capacity limit violations because it does not track the temporal usage of contacts. The linear volume model is illustrated in Figure 2 (right). As shown in the figure, the linear model incorrectly indicates that bundle $i$ (hatched rectangle) arriving at node C at time $t_{i}$ has immediate access to contact $\text{C}_{3}$. In contrast, the actual contact volume consumption (Figure 2, left) reveals that bundle $i$ cannot use contact $\text{C}_{3}$ immediately because bundle $a$ (striped rectangle) occupies the entire capacity at that time. When traditional CGR detects this conflict during forwarding through ETO, the transmission must be delayed.
@@
TABLE I: Initial Contact Plan.
Contact Sender Receiver Start time End Time
C1 A B $t_{1,1}$ $t_{2,1}$
C2 B C $t_{1,2}$ $t_{2,2}$
C3 C D $t_{1,3}$ $t_{2,3}$
C4 D E $t_{1,4}$ $t_{2,4}$
@@ III-B Buffer management
Node buffer management is crucial in DTNs because bundles may require extended storage before transmission. Buffer capacity constraints further complicate this challenge.
Traditional buffer management strategies react to problems after they occur. When a buffer is full, the system must immediately decide to reroute or drop bundles, as seen in reactive custody transfer [6]. An alternative approach proposed in [34] involves refusing custody to preserve node buffer space. However, these strategies only utilize information from neighboring nodes.
An alternative approach consists in implementing global buffer occupancy management during initial route planning. Unlike contact-level volume and capacity management, buffer occupation requires node-level management. Multiple bundles using different contacts can affect the same node’s buffer capacity. Consequently, this characteristic demands careful consideration when working with contact-graph network representations.
@@ IV Proposed global capacity and buffer management solution
Our solution builds upon bundle-dependent SR-CGR [31] (see Section II-D2). We address buffer and capacity constraints during the initial route-search phase. This proactive approach contrasts with existing reactive solutions that address issues after they occur.
Our approach guarantees that all discovered routes satisfy buffer and capacity constraints. The modifications also preserve all viable solutions, ensuring the discovery of optimal solution in the set of valid routes.
We achieve this by modifying the CP for both capacity and buffer constraints before route search begins. Additionally, we modify allowed edges between contacts during the route search to enforce buffer constraints. These modifications ensure CGR searches only within the space of valid solutions.
The system encodes the optimal route, which respects all buffer and capacity limits, into the bundle for network forwarding, following the standard source-routing CGR approach.
@@ IV-A Model-parameter modifications
  • First, we modify the CP through contact splitting for use in the route-search algorithm (i.e., Dijkstra search). The CP modifications fall into two categories:
    • Definitive modifications shared across the network.
    • Temporary and bundle-specific modifications applied locally during the route search.
  • Second, we restrict edges between contacts through specific rules. The route-search algorithm enforces these rules during local route computation for each bundle.
@@ IV-A1 Contact splitting to modify the CP
We introduce the contact splitting operation. Consider contact $\text{C}_{j}$ with data rate $R_{\text{C}_{j}}$ between nodes $M$ and $N$, active from $t_{1,j}$ to $t_{2,j}$. Contact splitting divides this contact into two new contacts $\text{C}_{j,1}$ and $\text{C}_{j,2}$ between $M$ and $N$: one from $t_{1,j}$ to $t_{e_{1}}$ and another from $t_{e_{2}}$ to $t_{2,j}$, where $t_{1,j}\leq t_{e_{1}}\leq t_{e_{2}}\leq t_{2,j}$ (see Figure 3). Both split contacts maintain the original data rate $R_{\text{C}_{j}}$.
When $t_{e_{2}}-t_{e_{1}}>0$, the operation erases part of the original contact. Additionally, if either time span equals zero, contact splitting results in contact shortening.
@@ IV-A2 Edge pruning
We define edge pruning as the act of forbidding edges between contacts. Edge pruning applies to any pair of contacts, whether original or split. In the example of Figure 2, one can choose to forbid the succesive use of contacts $\text{C}_{1}$ and $\text{C}_{2}$: pruning the edge between contacts $\text{C}_{1}$ and $\text{C}_{2}$ prevents their sequential use in any route.
Edge pruning occurs during the route search, specifically within the CRP. The implementation adds a new criterion to the contact exploration process (lines $8-12$ in Algorithm 1). When a contact fails this criterion, the algorithm:
  • Skips it during exploration.
  • Prevents it from becoming a successor to the currently selected contact $\text{C}_{s}$.
In essence, edge pruning enforces route constraints by preventing invalid contact sequences during the search process.
@@ IV-B1 Capacity management
This approach requires knowledge of each bundle’s size. The algorithm begins by searching a route for bundle $i$. Similarly to the linear-volume approach, the algorithm excludes contacts with insufficient volume for the bundle. Once the shortest path is selected for this bundle, the CP is modified as follows: We remove the part of the contact consumed by the bundle from the CP before performing the route search for subsequent bundles. The length of the contact to be erased is $t_{\text{Tx,}i}^{\text{C}_{j}}$, starting from the start time of the contact or the arrival time of the bundle at the contact’s sender node, whichever occurs later. The transmission time $t_{\text{Tx,}i}^{\text{C}_{j}}$ for bundle $i$ over contact Cj is:
$t_{\text{Tx,}i}^{\text{C}_{j}}=\frac{S_{i}}{R_{\text{C}_{j}}}.$ (1)
This contact splitting creates two new contacts (or one in case of shortening, see Section IV-A1) that replace the original one in the CP. We apply this operation to all contacts in the selected route. All subsequent route searches use this modified CP. These modifications are permanent and must propagate to all nodes performing route searches.
Example. In Figure 2, after the route search and the route selection for the initial bundle $a$ (striped rectangle), the CP becomes as represented in Table II. The new CP is then used for the route search of subsequent bundles. In this case, the part of contact $\text{C}_{3}$ between $t_{i}$ and $t_{i}+t_{\text{Tx,a}}^{\text{C}_{3}}$ is not an option for future bundle and no collision will therefore occur.
TABLE II: Modified Contact Plan after allocating route for bundle $a$.
Contact Sender Receiver Start time End Time
C1,1 A B $t_{1,1}$ $t_{a}$
C1,2 A B $t_{a}+t_{\text{Tx,}a}^{\text{C}_{1}}$ $t_{2,1}$
C2,1 B C $t_{1,2}+t_{\text{Tx,}a}^{\text{C}_{2}}$ $t_{2,2}$
C3,1 C D $t_{1,3}$ $t_{i}$
C3,2 C D $t_{i}+t_{\text{Tx,}a}^{\text{C}_{3}}$ $t_{2,3}$
C4 D E $t_{1,4}$ $t_{2,4}$
Complexity. The CGR complexity is directly related to the size of the CP. Contact splitting has two possible outcomes: either shortening (replacing with one contact) or splitting (replacing with two contacts). Each splitting operation increases the CP size by at most one: In a full split, one original contact becomes two new contacts. Shortening preserves the CP size by modifying the original contact’s duration. The CP size decreases by one when a shortened contact’s volume falls below the minimum bundle size. For bundle $i$, the number of split contacts equals the number of hops in its route $N_{\text{hops},i}$. In a time window $w$ with $N_{b,w}$ routed bundles, the maximum number of added contacts to the CP is therefore:
$\Delta_{\text{CP},w}^{+}=\sum_{i=1}^{N_{b,w}}{N_{\text{hops},i}}.$ (2)
Our implementation removes contacts (split or original) from the CP upon expiration. Let $N_{\text{exp},w}$ represent the number of expired contacts in time window $w$. The resulting reduction in CP size is simply $\Delta_{\text{CP},w}^{-}=N_{\text{exp},w}$.
Finally, the net change in the CP size during window $w$ is:
$\Delta_{\text{CP},w}=\Delta_{\text{CP},w}^{+}-\Delta_{\text{CP},w}^{-}.$ (3)
Section VII shows our algorithm’s impact on the CP size, and thus on the CGR complexity, through practical examples.
@@ IV-B2 Buffer management
We introduce forecast buffer tables as a key component of our solution. These tables track buffer occupancy status for each node over time. Similarly to the CP, all route-computing entities share these tables. These tables then guide temporary modifications to both the CP and allowed edges.
Each node has a maximum buffer capacity $B_{\text{max}}$, with overflow occurring beyond this limit. Our buffer management solution applies temporary and bundle-specific CP modifications. Edge pruning follows the same temporary and bundle-specific pattern. These modifications remain local to the computing node and apply only during individual route searches.
Forecast buffer tables. Once a route is selected, the buffer tables are updated accordingly and shared to all entities performing a route search. In other words, once a route is validated, one checks when the bundle will reach a certain node and how long will it stay in its buffer according to the validated route.
@@
Algorithm to respect the buffer constrains. The proposed algorithm enforces buffer-limit constraints through the following approach: Using bundle size and forecast buffer tables, the algorithm verifies buffer availability at each node throughout the route. The algorithm excludes routes that would cause buffer overflow at any node. First, it modifies the CP temporarily through contact splitting. Next, it prunes specific edges during the route search. This algorithm can be formulated with the four following main steps.
@@
Example. Some problematic contacts for the potential overflow $o_{1}$ described in the previous example are shown in Figure 6. These contacts have node $Z_{1}$ as receiver (regardless of their sender node $X$), start before $t_{2,o_{1}}$, and end after $t_{1,o_{1}}$. They are problematic because their duration intersects the period of overflow $o_{1}$. The problematic portions of these contacts are horizontally striped in Figure 6. They may forward bundle b to $Z_{1}$ during the potential overflow $o_{1}$ duration. The parts of these contacts that intersect the overflow duration are to be temporarily removed using contact splitting. The same applies to all potential overflows at all nodes.
@@
Example. After step 2, there is no contact having $Z_{1}$ as receiver that spans the duration of $o_{1}$ in Figure 6. Nevertheless, there are contacts having $Z_{1}$ as receiver and ending before $t_{1,o_{1}}$. Using one of these contacts (e.g., $\text{C}_{i}$), a bundle arrives to $Z_{1}$ before the potential overflow $o_{1}$. However, this bundle must leave $Z_{1}$ before the beginning of overflow $o_{1}$ to avoid the potential problem. The bundle leaves node $Z_{1}$ through a contact succeeding the previous one ($\text{C}_{i}$) in a route. The succeeding contact (e.g., $\text{C}_{j}$) has $Z_{1}$ as sender, regardless of the receiver. As a result, the succeeding contact $\text{C}_{j}$ should start before the potential overflow $o_{1}$ to avoid it. Otherwise, the succession of $\text{C}_{i}$ and $\text{C}_{j}$ in a route causes the overflow and is therefore problematic.
@@
Step 4: Finally, the last step is to forbid edges between contacts yielding problematic contact successions. In other words, the edge with the contacts starting too late are pruned. This edge pruning is performed during the route search, in the CRP: lines $8-12$ in Algorithm 1. Note that additional inputs were required for the CRP to make our edge pruning possible: line $2$ in Algorithm 1. In addition to its original inputs, the CRP takes the bundle $b$ and the overflows start and end times $T_{1,s}$ and $T_{2,s}$ at the receiver of the selected contact $\text{C}_{s}$, respectively.
Example. In the example of Figure 6, during the route search, if a contact $\text{C}_{s}$ forwards the bundle to node $Z_{1}$ before $o_{1}$, the edge between $\text{C}_{s}$ and a contact $\text{C}_{i}$ from $Z_{1}$ to a third node is only accepted if $\text{C}_{i}$ starts before $t_{1,o_{1}}$. The same applies to all overflows at all nodes. More precisely, let us consider, in a CRP iteration, that $\text{C}_{s}$ is the selected contact and $\text{C}_{i}$ is another contact in the CP. Also, let us consider that the receiver of $\text{C}_{s}$ is node $Z_{1}$, having one potential overflow $o_{1}$. To consider $\text{C}_{i}$ as explored or a candidate successor for $\text{C}_{s}$, it must pass all the checks in the original CRP (lines $5-7$ in Algorithm 1). In addition to these checks, we add the following condition: If $t_{\text{arr},s}$ the bundle arrival time to the receiver of $\text{C}_{s}$ is before $t_{1,o_{1}}$ ($t_{\text{arr},s}<t_{1,o_{1}}$), then the start time of $\text{C}_{i}$ must also be before $t_{1,o_{1}}$ ($t_{1,i}<t_{1,o_{1}}$). Otherwise, $\text{C}_{i}$ is ignored in the CRP and it is not explored. If this condition is not met, $\text{C}_{i}$ cannot be a candidate successor for $\text{C}_{s}$. In a more general scenario, where multiple potential overflow periods exist at a node buffer, a bundle arriving at the node via contact $\text{C}_{s}$ between two overflow periods should depart from the node using contact $\text{C}_{i}$ before the next overflow period begins (lines $8-12$ in Algorithm 1).
@@
As described by the above steps, only invalid routes are removed by the modifications of the proposed contact capacity and buffer management solutions. The proposed modifications before and during the route search do not omit any feasible solution. Hence, the subsequent use of CGR gives the optimal route in the subset of valid solutions.
Complexity. As already mentioned, the CGR complexity directly depends on the CP size. Buffer management temporarily increases the CP size when routing bundle $b$, based on the bundle’s size and current buffer occupancy. The size increase depends on two factors: the number of overflowed nodes $|\mathcal{Z}_{\text{b}}|$ and the contacts intercepting overflows at these nodes. Each node $Z_{k}$ can experience multiple overflows, denoted as $|\mathcal{O}_{Z_{k}}|$. When a contact intersects an overflow period, it splits into at most two new contacts, increasing the CP size by one. This increase remains temporary and applies only during bundle $b$’s routing phase. After routing completes, the system reverts to the permanent CP. At most, the number of temporarily added contacts to the CP for routing bundle $b$ with buffer constraints is:
$\Delta_{\text{CP},b}^{+}=\sum_{k=1}^{|\mathcal{Z}_{\text{b}}|}\sum_{l=1}^{|\mathcal{O}_{Z_{k}}|}{N_{\text{c}_{l,k}}},$ (4)
where $N_{\text{c}_{l,k}}$ represents the number of contacts having node $Z_{k}\in\mathcal{Z}_{\text{b}}$ as destination and intercepting overflow $o_{l}\in\mathcal{O}_{Z_{k}}$. Section VII shows the practical impact of this management solution on the CP size.